OpenStudy (wesdg1978):

lim x-->-pie sqrt(x+8) cos(x+pie) How do you find the limit? I don't have a clue where to begin.

OpenStudy (anonymous):

I think you are thinking way too hard.\[\cos(-\pi + \pi) = \cos(0) = 1\]\[\sqrt{-\pi+8}=\sqrt{8-\pi}\]\[1*\sqrt{8-\pi}=\sqrt{8-\pi}\]

OpenStudy (wesdg1978):

I think you're right. Looking at how easy that looks the way you wrote, I feel kinda stupid.

OpenStudy (wesdg1978):

The professor said that we have to "earn" our answers and very few will be able to be answered by substitution. So everyone I see, I think has some special trick to it.

OpenStudy (wesdg1978):

Thanks for your help!

OpenStudy (anonymous):

Anytime :)